### v1

# Superconducting Diode Effect in Weak Localization Regime

Abstract
We study a dirty two-dimensional superconductor with Rashba spin-orbit coupling and in-plane Zeeman fields described by the nonlinear sigma model that includes short-range electron-electron interactions from the Coulomb and Cooper channels. The renormalized Ginzburg-Landau theory, which includes the weak localization effects at the one-loop level, is constructed by using the Keldysh functional formalism. It is shown that the tricritical point appears in the phase diagram. The superconducting diode quality factor increases divergently as the system approaches the tricritical point. Near the superconducting phase transition lines, the absolute value of the diode quality factor decreases due to the cooperation of localization and interactions. The normal conductivity of the resistive state, in which the superconducting state is destroyed by the critical current, is calculated, and localization behaviors are demonstrated.
Introduction: The phenomenon in which the critical currents of superconductors become nonreciprocal, known as the superconducting diode (SD) effect [1, 2], has attracted much attention. Theoretical studies have been carried out for various types of noncentrosymmetric superconductors [3, 4, 5, 6, 7, 8, 9]. Not only clean systems but also systems with uncorrelated impurities have been studied in detail at the mean-field level [10]. If electron systems are disordered enough, they are well described by the nonlinear sigma model (NLSM) [11, 12, 13, 14], and dirty superconductors can also be described by the NLSM [15, 16, 17, 18]. Recently, an NLSM for a dirty superconductor with spin-orbit coupling (SOC) and a Zeeman field was derived [19] and used to investigate the SD effect in Josephson junctions at the mean-field level [20]. However, the Anderson localization effect, a correction to the mean-field theory in dirty systems, has not yet been highlighted in studies of the SD effect.
The role of Anderson localization in superconductors is an important topic of ongoing debate. Without considering strong mesoscopic correlations and electron-electron (e-e) interactions, the properties of dirty $s$-wave superconductors at zero magnetic field remain unchanged with respect to the strength of non-magnetic disorder in the weak localization (WL) regime, known as the Anderson theorem [21]. If the Coulomb interaction is concerned, the transition temperature for the $s$-wave pairing states is suppressed in the WL regime [22, 23, 24]. The opposite effect can occur in the strong localization regime by the multifractality effect [25, 26, 27, 28]. It was also shown that, in the WL regime, the Coulomb interaction suppresses the critical current of conventional dirty superconductors at zero magnetic field and low temperatures [29, 30].
In this Letter, we study the interplay of the e-e interaction and disorder on the SD effect and clarify the WL correction to conductivity when the superconducting (SC) state is broken by the critical current. We work with the NLSM for a dirty two-dimensional (2D) noncentrosymmetric superconductor, including the interactions from the Coulomb and Cooper channels in the short-range limit, and establish a renormalized Ginzburg-Landau (GL) theory at the one-loop level. While many studies [28, 31, 32, 33] have been conducted in the replica formalism, we formulate the theory with the Keldysh functional formalism [14, 15, 17, 34].
Model: We consider a dirty 2D superconductor aligned in the $xy$-plane with the Rashba SOC and a Zeeman field in the $x$-direction. In the Keldysh functional formalism, the partition function is given as $Z=\int D[Q,\Delta]\exp\left({iS[Q,\Delta]}\right)$ with [19]
$S[\Delta,Q]$ $=S_{\rm HS}[\Delta]+S_{\rm N}[\Delta,Q]+S_{\rm T}[Q],$ (1)
where $S_{\rm HS}$ is responsible for the Hubbard-Stratonovich transformation associated with the SC state, $S_{\rm N}$ is the conventional non-interacting part, and $S_{\rm T}$ is the topological part. Their explicit forms are shown in Supplemental material [35]. A $\times 8$ 8 matrix field $Q$ satisfies the conditions $Q^{2}=1$ and $\mathrm{Tr}\,\{Q\}=0$. Let $\gamma_{\mu},~\tau_{\mu}$, and $\sigma_{\mu}~(\mu=0,\dots,4)$ be the Pauli matrices in the Keldysh, Nambu, and spin spaces, respectively. In particular, we write $\gamma_{\rm c}=\gamma_{0}$ and $\gamma_{\rm q}=\gamma_{1}$ and often abbreviate $\gamma_{0}$ for simplicity. The order parameter of superconductivity, that is, the SC gap, is then described by the matrix $\gamma_{\rm c}\check{\Delta}_{\rm c}+\gamma_{\rm q}\check{\Delta}_{\rm q}$, where $\check{\Delta}_{\rm c(q)}=\tau_{+}\Delta_{\rm c(q)}+\tau_{-}\Delta^{*}_{\rm c(q)}$ with $\tau_{\pm}=(\tau_{1}\pm i\tau_{2})/2$ and $\Delta_{\rm c(q)}$ is the classical (quantum) component of the SC gap. The check symbol $\check{(\cdot)}$ highlights that the objects are $\times 4$ 4 matrices. The terms in the form $\gamma_{\mu}\check{A}$ stand for the expression $\gamma_{\mu}\otimes\check{A}$, where $\otimes$ is a tensor product. We assume that the SC state is a helical state modulating in the $y$-direction and time-independent. In this case, the Cooper pair momentum corresponds to a uniform gauge field along the $y$-axis $A_{y}$, and we set $A_{x}=0$ from now on.
In addition to Eq. (1), we take into account interaction terms of the Cooper and Coulomb channels. The former can be obtained directly by integrating the spatial fluctuation of the gap function from Eq. (1) [18, 28]. The procedure results in two terms in action. The first term $\bar{S}_{\rm HS}$ has the same form as $S_{\rm HS}$ [Eq. (S.1) in Supplemental Material [35]], but the spatial modulation in the gap functions is removed. Thus, in the following, $\Delta_{\rm c(q)}$ denotes the spatially uniform gap function. The second term is
$S^{\rm(c)}_{\rm int}[Q]=\tilde{\Gamma}_{\rm c}\!\!\sum\limits_{i=1,2}\!\int_{t,\bm{r}}\!\!\!\mathrm{tr}\!\left\{{\gamma_{\rm c}\tau_{i}\delta Q_{tt}(\bm{r})}\right\}\!\mathrm{tr}\!\left\{{\gamma_{\rm q}\tau_{i}\delta Q_{tt}(\bm{r})}\right\}\!,$ (2)
[…]
$S^{\rm(s)}_{\rm int}[Q]=\tilde{\Gamma}_{\rm s}\int_{t,\bm{r}}\mathrm{tr}\left\{{\gamma_{\rm c}\tau_{0}Q_{tt}(\bm{r})}\right\}\mathrm{tr}\left\{{\gamma_{\rm q}\tau_{0}Q_{tt}(\bm{r})}\right\},$ (3)
where $\tilde{\Gamma}_{\rm s}=-\pi^{2}\nu\Gamma_{\rm s}/16$ and $\Gamma_{\rm s}=1$ in the short-range limit. Thus, the total action in the partition function $Z$ is given by $S_{\rm tot}=\bar{S}_{\rm HS}+S_{\rm N}+S_{\rm T}+S^{\rm(c)}_{\rm int}+S^{\rm(s)}_{\rm int}$.
Effective action: In the ordinary mean-field theory, the system is described by saddle points of action, which are obtained by searching the matrix field $Q=\bar{Q}$ that optimizes the action under the constraints $\bar{Q}^{2}=1$ and $\mathrm{Tr}\,\{\bar{Q}\}=0$. Going beyond the mean-field approximation can be done by integrating out the fluctuations around the saddle points, resulting in the effective action of the gap function. In this work, we consider the Gaussian fluctuations around a SC saddle point $\bar{Q}_{\rm s}$. Since $\bar{Q}_{\rm s}$ must be on the same manifold as a normal saddle point $\bar{Q}_{\rm N}$, there is a matrix $U$ such that $\bar{Q}_{\rm s}=U\bar{Q}_{\rm N}U^{-1}$. The saddle point configuration may be taken as the fermionic action formulated by Keldysh path integration [14]. That is $\bar{Q}=U_{\rm K}\Lambda U_{\rm K}$, where $\Lambda=\mathrm{diag}\,\{\check{\Lambda}_{\rm R},\check{\Lambda}_{\rm A}\}$ and $\rm R(A)$ denotes the retarded (advanced) component. In the energy representation, $U_{\rm K}$ is
$U_{\mathrm{K}\epsilon}$ $=\begin{pmatrix}\tau_{0}&F_{\epsilon}\tau_{0}\\ 0&-\tau_{0}\end{pmatrix},$ (4)
where distribution in thermal equilibrium leads to $F_{\epsilon}=\tanh\epsilon/(2T)$ with energy $\epsilon$ and temperature $T$. Using the above equations, we can rewrite $\bar{Q}_{\rm N(s)}=U_{\rm K}\Lambda_{\rm N(s)}U_{\rm K}$, where $\Lambda_{\rm N(s)}=\mathrm{diag}\,\{\check{\Lambda}_{\rm N(s),R},\check{\Lambda}_{\rm N(s),A}\}$. The form of $\Lambda_{\rm N}$ in the energy representation is $\Lambda_{\mathrm{N}\epsilon}=\mathrm{diag}\left\{{\tau_{3},-\tau_{3}}\right\}$ [14, 15]. On the other hand, $\Lambda_{\rm s}$ is generally a function of the spectral angle $\theta$, which will be specified later. Note that $U$ is also a function of $\theta$, just as $\Lambda_{\rm s}$ is.
Next, we perform the transformation $Q=U\tilde{Q}U^{-1}$ and then the transformation $\tilde{Q}=\exp\left({W/2}\right)\Lambda_{\rm N}\exp\left({-W/2}\right)$. Thus, $Q$ can be expanded to the quadratic order of $W$ as $Q=\Lambda_{\rm N}+\frac{1}{2}\,[W,\Lambda_{\rm N}]+\frac{1}{8}\,[W,[W,\Lambda_{\rm N}]]$. Expanding the total action to the quadratic order of $W$ results in $S_{\mathrm{tot}}[\Delta,Q]\rightarrow S_{\mathrm{tot}}[\Delta,\theta,W]$, where
$S_{\mathrm{tot}}[\Delta,\theta,W]$ $=S_{0}[W]+S_{\rm MF}[\theta,\Delta]+\tilde{S}_{\rm int}[\theta,W]$
$\quad+S_{\rm I}[\theta,W]+S_{\rm II}[\theta,\Delta,W].$ (5)
Here, $S_{0}[W]$ is the action of the free cooperon and diffuson modes, which are determined from the condition $\{W,\Lambda_{\rm N}\}=0$. The explicit expressions for each term except the last term in Eq. (5) are listed in [35]. The last term $S_{\rm II}$ contains the terms in the form $\sim\mathrm{Tr}\,\{(\dots)WW\}$ and $\sim\mathrm{Tr}\,\{W(\dots)W\}$, which do not make any contribution in the first order of $\Delta_{\rm q}$ and thus are neglected.
Now, we can find the effective action of the gap function, $S_{\rm eff}$. Regarding the terms other than $S_{0}[W]$ as perturbative terms, $S_{\rm eff}$ is given as $\exp\left({iS_{\rm eff}}\right)=\braket{\exp\left({i\left[{S_{\rm MF}+S_{\rm int}+S_{\rm I}}\right]}\right)}_{0}$, where $\braket{\cdots}_{0}$ denotes the average with respect to $S_{0}$. Performing the cumulant expansion, at the one-loop level, the effective action is obtained from the following equation:
$S_{\rm eff}=\braket{S_{\rm MF}+\tilde{S}_{\rm int}}_{0}+\frac{i}{2}\braket{S^{2}_{\rm I}}_{0}.$ (6)
SC saddle point: The SC saddle points $\bar{Q}_{\rm s}$ are usually obtained by numerical techniques. However, since we aim to integrate out the fluctuations to obtain an effective action of the SC gap function, saddle points can be approximated so that all the fluctuations can be considered. Although taking into account all fluctuations is a difficult task, a method can give the familiar GL action for conventional dirty superconductors by calculations at the tree level [16, 17]. Here, we choose the time-independent SC saddle point $\bar{Q}_{\rm s}$ that includes the Zeeman coupling but excludes the interaction terms, the SOC, and the Cooper pair momentum. The saddle point $\bar{Q}_{\rm s}$ is obtained by solving the equation $[\check{E}_{\epsilon}+\check{\Delta}_{\rm c}+\gamma_{\rm q}\check{\Delta}_{\rm q},\bar{Q}_{\mathrm{s},\epsilon\epsilon}]=0$, where $\check{E}_{\epsilon}=\tau_{3}(\epsilon+h\sigma_{3})$ and $h$ is a Zeeman field. The equation can be easily solved if $\Delta_{\rm q}=0$. However, in the absence of the quantum part $\Delta_{\rm q}$, the solution cannot include any effect of the terms other than the term associated with $\check{E}$ on the effective action. We overcome this problem by instead using the solution of the following equation:
$[\check{E}_{\epsilon}+\check{\Delta}_{\mathrm{R(A)},\epsilon}$ $,\check{\Lambda}_{\mathrm{sR(A)},\epsilon\epsilon}]=0,$ (7)
$\check{\Delta}_{\mathrm{R(A)},\epsilon}=$ $\check{\Delta}_{\rm c}\pm F_{\epsilon}\check{\Delta}_{\rm q}.$ (8)
Solving Eq. (7), we obtain the saddle point $\bar{Q}_{\rm s}=U_{\rm K}\Lambda_{\rm s}U_{\rm K}$ with $\Lambda_{\rm s}=\mathrm{diag}\,\{\check{\Lambda}_{\rm sR},\check{\Lambda}_{\rm sA}\}$, which can be written in the form $\check{\Lambda}_{\rm sR(A)}=\check{U}_{\rm sR(A)}\tau_{3}\check{U}^{-1}_{\rm sR(A)}$. By some calculations, we obtain $\check{U}_{\rm sR(A)}=1-\tau_{1}\hat{X}_{\rm R(A)}$, where
$\hat{X}_{\rm R}=\mathrm{diag}\,$$\Big{\{}\tanh\frac{\theta_{\mathrm{R},\epsilon+}}{2},\tanh\frac{\theta_{\mathrm{R},\epsilon-}}{2}\Big{\}},$ (9)
$\tanh\theta_{\mathrm{R},\epsilon\pm}$ $=\mathrm{sgn}(\epsilon_{\pm})\Delta_{\mathrm{R},\epsilon}/|\epsilon_{\pm}\!+\!i0|,$ (10)
and $\epsilon_{\pm}=\epsilon\pm h$. $\hat{X}_{\rm A}$ has the same form as $\hat{X}_{\rm R}$ but $\theta_{\mathrm{R},\epsilon\pm}$ is replaced by $\theta_{\mathrm{A},\epsilon\pm}$, which satisfies $\tanh\theta_{\mathrm{A},\epsilon\pm}=\Delta_{\mathrm{A},\epsilon}/|\epsilon_{\pm}\!-\!i0|$. This is indeed a possible form of the spectral angle mentioned below Eq. (4). Defining $U\equiv U_{\rm K}U_{\rm s}$ and $U_{\rm s}\equiv\mathrm{diag}\left\{{\check{U}_{\rm sR},\check{U}_{\rm sA}}\right\}$, we can rewritte $\bar{Q}_{\rm s}$ as $\bar{Q}_{\rm s}=U\Lambda_{\rm N}U^{-1}$.
The solution of Eq. (7) can be used to obtain the effective action including the contributions of the SOC, Cooper-pair momentum, and interactions. However, we will go to a more accurate approximation by allowing the spectral angles $\theta_{\mathrm{R(A)},\epsilon\pm}$ to be determined by optimizing the total effective action with respect to them. The outcome is the so-called modified Usadel equation, first proposed in Ref. [28] by the replica formalism. To derive a similar equation in the Keldysh functional formalism, we have to take into account the quantum part of the spectral angle. Otherwise, for the same reason as described in the previous paragraph, the effective action becomes trivial. The problem is how the quantum part of the spectral angles should enter $\Lambda_{\rm s}$. By observing Eqs. (8) and (10), we arrive at the ansatz that the generalized spectral angles should have the form like Eq. (8) for the gap function, and the spectral angles in Eq. (9) should be replaced as follows:
$\theta_{\mathrm{R},\epsilon\pm}\!$ $\rightarrow\theta_{\mathrm{R},\epsilon\pm}=\theta^{\rm c}_{\mathrm{R},\epsilon\pm}\!\!+F_{\epsilon}\theta^{\rm q}_{\mathrm{R},\epsilon\pm},$ (11)
$\theta_{\mathrm{A},\epsilon\pm}\!$ $\rightarrow\theta_{\mathrm{A},\epsilon\pm}=\theta^{\rm c}_{\mathrm{A},\epsilon\pm}\!\!-F_{\epsilon}\theta^{\rm q}_{\mathrm{A},\epsilon\pm},$ (12)
where $\theta^{\rm c}_{\mathrm{R(A)},\epsilon\pm}$ and $\theta^{\rm q}_{\mathrm{R(A)},\epsilon\pm}$ represent the classical and quantum parts of the generalized spectral angles, respectively. In the above equation, $F_{\epsilon}$ plays an essential role. The effective action would become trivial if there were no term that couples $F_{\epsilon}$ to $\theta^{\rm q}_{\mathrm{R(A)},\epsilon\pm}$. Hereafter, we adopt the SC saddle point $\bar{Q}_{\rm s}$ that is determined by Eq. (9) with Eqs. (11) and (12). The quantum parts will be set to zero after performing the first derivatives with respect to them.
Modified Usadel equation: Following the plan described in the previous paragraph, we determine the form of $\theta^{\rm c}_{\mathrm{R(A)},\epsilon\pm}$ by optimizing the effective action $S_{\rm eff}$ with respect to each spectral angle. We work within GL theory by expanding $\theta_{\mathrm{R(A)},\epsilon\pm}$ as $\theta^{\rm c}_{\mathrm{R(A)},\epsilon\pm}=\Delta_{\rm c}f^{\rm R(A)}_{1\epsilon\pm}+\Delta^{3}_{\rm c}f^{\rm R(A)}_{3\epsilon\pm}$ and $\theta^{\rm q}_{\mathrm{R(A)},\epsilon\pm}=\theta^{\rm R(A)q}_{1\epsilon\pm}+\theta^{\rm R(A)q}_{3\epsilon\pm}$. In this way, $f^{\rm R(A)}_{j\epsilon\pm}$ is determined by the following equation:
$[\delta S_{\rm{eff}}/\delta\theta^{\mu\mathrm{q}}_{j\epsilon\pm}]_{\theta^{\mathrm{q}}\rightarrow 0}=0,$ (13)
where $\mu=\rm R,A$ and $j=1$ or 333. In the following, $f^{\rm R(A)}_{j\epsilon\pm}$ is referred to as spectral function. In the calculation of Eq. (13), the retarded and advanced parts are not independent. Otherwise, irregular divergences appear in the left-hand side of Eq. (13). This argument is explained in detail in Supplemental Material [35]. We find that all irregular terms vanish if the following condition holds:
$f^{\rm R}_{j\epsilon\pm}=-f^{\rm A}_{j-\epsilon\mp},\quad\theta^{\rm Rq}_{j\epsilon\pm}=-\theta^{\rm Aq}_{j-\epsilon\mp}.$ (14)
Using Eqs. (11)--(14) and performing analytic continuation, we obtain the modified Usadel equations at the first and third orders of $\Delta_{\rm c}$ as follows:
$\mathcal{S}_{\mathrm{kernel},j}[f_{1n\pm},f_{3n\pm}]=\mathcal{S}_{\mathrm{source},j}[f_{1n\pm},f_{3n\pm}],$ (15)
where $f_{jn\pm}=if^{\rm R}_{ji\omega_{n}\pm}$, $\omega_{n\pm}=\omega_{n}\pm ih$, $\omega_{n}=(2n+1)\pi T~(n\in\mathbb{Z})$, and $\mathcal{S}_{\mathrm{source},1}=1$. The others are shown in Supplemental Material [35] to avoid describing the lengthy equations, and we find that the Cooper interaction terms are not involved in the equation determining $f_{1n\pm}$. The equations are solved numerically by the Padé decomposition method [36], which can achieve fast convergence when calculating the Matsubara summation.
Equation of state: The gap equation is obtained from the stationary condition $[\partial S/\partial\Delta_{\rm q}]_{\Delta_{\rm q}\rightarrow 0}=0$. In GL theory, the equation has the form $\alpha_{2}\Delta_{\rm c}+\alpha_{4}\Delta_{\rm c}^{3}=0$, where
$\alpha_{2}$ $=4\,\Gamma^{-1}_{\rm c}\!-4\pi T\!\sum_{n>0}\sum_{\lambda=\pm}\Big{[}f_{1n\lambda}\!+\!\frac{1}{\omega_{n}\!+\!i\lambda h}\Big{]}$
$\quad-2\int_{0}^{\omega_{\rm D}}\frac{d\epsilon}{\epsilon}\Big{[}\tanh\left({\frac{\epsilon_{+}}{2T}}\right)+\tanh\left({\frac{\epsilon_{-}}{2T}}\right)\Big{]},$ (16)
$\alpha_{4}$ $=-4\pi T\!\sum_{n>0}\sum_{\lambda=\pm}\Big{(}f_{3n\lambda}-\!\frac{1}{6}[f_{1n\lambda}]^{3}\Big{)}.$ (17)
Note that $f_{jn\lambda}$ is derived from Eq. (15), and $\alpha_{2}$ is regularized by introducing the ultraviolet cutoff, which is the Debye energy $\omega_{\rm D}$, to the energy integration. The form of the gap equation in the GL level suggests that the free energy and the supercurrent are given as $f_{\mathrm{GL}}=\frac{1}{2}\alpha_{2}\Delta^{2}_{\rm c}+\frac{1}{4}\alpha_{4}\Delta^{4}_{\rm c}$ and $I_{\rm s}=\partial f_{\mathrm{GL}}/\partial A_{y}$, respectively. The helical SC state for any supercurrent $I_{s}$ is then determined by the latter equation. In particular, the second-order SC phase transition line in the absence of the supercurrent is calculated based on the conditions $\alpha_{2}=0$ and $\partial\alpha_{2}/\partial A_{y}=0$. From now on, let $A_{0}$ be the $A_{y}$ field in the zero-current SC states. We denote the transition temperature at a magnetic field $h$ as $T_{\rm c}(h)$, which is denoted as $T_{\rm c0}(h)$ when we neglect the e-e interaction. At $h=0$, we describe $T_{\rm c0}(0)=T_{\rm c0}$ and $T_{\rm c}(0)=T_{\rm c}$, for simplicity. For later use, we define the reduced temperature $\tau_{\rm r}$ as $\tau_{\rm r}=T/T_{\rm c0}$ or $\tau_{\rm r}=T/T_{\rm c}$.
WL conductivity: When the current reaches its critical value, the SC state changes to a resistive state. Here, we study the WL correction to the conductivity, $\sigma_{\rm WL}$ [37, 38, 39]. For simplicity, we neglect the e-e interactions. We find that it is convenient to work with the alternative normal saddle point $\bar{Q}^{\prime}_{\rm N}=U^{\prime}\Lambda^{\prime}_{\rm N}U^{\prime-1}$, where $\Lambda^{\prime}_{\rm N}=\mathrm{diag}\,\{\gamma_{3},\gamma_{3}\}$ and $U^{\prime}=\mathrm{diag}\,\{\check{U}_{\rm K},\check{U}^{\mathbb{T}}_{\rm K}\}$. Here, $U^{\mathbb{T}}$ denotes the transpose of a matrix and the interchange of the time of $U$, which means that $U^{\mathbb{T}}_{t_{1}t_{2}}=U^{\rm t}_{t_{2}t_{1}}$ with $U^{\rm t}_{t_{2}t_{1}}$ being the ordinary transpose in matrix elements of $U_{t_{2}t_{1}}$. The matrix $\check{U}_{\rm K}$ has the same form as Eq. (4), but $\tau_{0}$ is replaced by $\sigma_{0}$.
This saddle point comes from the time-reversal symmetry condition of the matrix field $Q$. To the best of our knowledge, this saddle point was first introduced in Ref. 40 to calculate the WL conductivity $\sigma_{\rm WL}$ in the spinless case within the Keldysh functional formalism. Here, we generalize their formalism to include spin degrees of freedom. We explain the procedure in [35], and the equation for $\sigma_{\rm WL}$ is consistent with existing results [41, 42, 44, 43].
Numerical results: In the numerical calculations, we fix the parameters, $\Gamma_{\rm c}=0.2$, $\omega_{\rm D}/T_{\rm c0}=0.1309$, and Fermi energy $E_{\rm F}/T_{\rm c0}=p_{\rm F}^{2}/(2mT_{\rm c0})=10^{3}$, and consider the following two sets of parameters: (I) $\alpha/p_{\rm F}=0.01,~\tau T_{\rm c0}=1.5\times 10^{-2}$ and (II) $\alpha/p_{\rm F}=0.002,~\tau T_{\rm c0}=7.5\times 10^{-3}$, where $\alpha$ is the Rashba SOC constant and $\tau$ is the impurity scattering time. Note that the values of $\tau T_{\rm c0}$ in cases (I) and (II) correspond to the characteristic lengths of $l_{\rm F}p_{\rm F}=30$ and 151515, respectively, which are consistent with the WL regime. The characteristic energies of SOC and disorder are represented by $\epsilon_{\rm SOC}=\alpha v_{\rm F}$ and $\epsilon_{\rm D}=1/\tau$, respectively. In case (I), the SOC is comparable with disorder as $\epsilon_{\rm SOC}/\epsilon_{\rm D}=0.6$, while case (II) is in a weak SOC regime as $\epsilon_{\rm SOC}/\epsilon_{\rm D}=0.06$. We show the numerical results for the comparable and weak SOC in Figs. 1 and 2, respectively.
Figure 1: The diode quality factor $\eta$ and the WL correction to conductivity $\sigma_{\rm WL}$ in the intermediate SOC regime ($\alpha/p_{\rm F}=0.01,~\tau T_{\rm c0}=1.5\times 10^{-2}$). The solid (dashed) lines show the results in the interacting (free) case with (without) the e-e interaction. Panels (a) and (c) show the color maps of $\sigma_{\rm WL}$ and $\eta$, respectively. The solid and dashed lines show the second-order superconducting transition line. The bold circles in (a) represent the tricritical points. Panels (b) and (d) show $\sigma_{\rm WL}$ and $\eta$ near the SC transition line, respectively. In these figures, we vary the reduced temperature $\tau_{\rm r}$ and magnetic field $h$ so that the temperature is $T=0.99T_{\rm c}(h)$.
We first show the properties of zero-current SC states. As shown in Figs. 1a and 2a, the e-e interactions suppress the transition temperatures and critical magnetic fields. The gauge field $A_{0}$ corresponding to the momentum of Cooper pairs is also suppressed when the interactions are incorporated [35]. Since there is no contribution from the Cooper channel in the equation determining the second-order SC transition line, the suppression of the transition temperature is attributed to the Coulomb channel. The zero-field transition temperatures are $T_{\rm c}=0.8793T_{\rm c0}$ and $T_{\rm c}=0.6457T_{\rm c0}$ in Figs. 1 and 2, respectively. We have confirmed that $T_{\rm c}$ decreases as the impurity scattering time $\tau$ decreases, while it increases toward $T_{\rm c0}$ as the SOC $\alpha$ increases. The transition lines depicted in Figs. 1a and 2a show the second-order phase transition, and they terminate at the tricritical points at high magnetic fields, indicating the first-order transition, at which $\alpha_{4}$ becomes negative. It is well known that the first-order SC transition occurs in the high field region due to the Pauli paramagnetic effects when we adopt the mean-field theory for spatially uniform $s$-wave states [45]. At magnetic fields higher than the tricritical point, the second-order transition from the normal state to the non-uniform stripe SC state is expected [46]. Since the Cooper channel makes a non-zero contribution to $\alpha_{4}$, both Coulomb and Cooper channels affect the tricritical points.
Figure 2: The WL correction to conductivity $\sigma_{\rm WL}$ in the weak SOC regime ($\alpha/p_{\rm F}=0.002,~\tau T_{\rm c0}=7.5\times 10^{-3}$). Panel (a) shows the color map of $\sigma_{\rm WL}$, which is plotted in the same manner as Fig. 1a. Panel (b) shows $\sigma_{\rm WL}$ near the superconducting transition line in the same manner as Fig. 1b.
Next, let us focus on the quality factor $\eta$ of the SD effect. The diode quality factor $\eta$ is defined as $\eta=(1-\xi)/(1+\xi)$, where $\xi=|I_{\rm sc-}/I_{\rm sc+}|$ and $I_{\rm sc\pm}$ is the critical current in the ±$\pm y$-direction. Because $\eta$ is tiny in the weak SOC regime, we calculate the SDE in the intermediate SOC regime and show the results of $\eta$ in Figs. 1c and 1d. As shown in Fig. 1c, the diode quality factor increases as the system approaches the tricritical point. To show the behaviors of the SDE near the transition line, we plot in Fig. 1d the diode quality factor near the transition line as a function of the reduced temperature. The absolute values of the diode quality factor increase divergently as the tricritical point is approached. We see qualitatively the same behaviors in the absence and presence of the e-e interaction, while the tricritical point is shifted. Although the Coulomb interaction significantly reduces the zero-field transition temperature, the diode quality factor almost coincides in both cases at high reduced temperatures. This behavior is deviated as the tricritical points are approached. Because the tricritical point is shifted by interactions, the absolute values of the diode quality factor in the presence of interactions become smaller than those in the case without interactions.
Finally, we discuss the WL correction to the conductivity $\sigma_{\rm WL}$. The correction is normalized with the Drude conductivity $\sigma_{\rm D}=2\nu D$. As shown in Fig. 1a, interestingly, $\sigma_{\rm WL}$ in the intermediate SOC regime decreases as the temperature and magnetic field increase, indicating stronger localization effects. By examining the plot near the transition lines [Fig. 1b], we conclude that the e-e interaction can suppress the WL correction. However, as shown in Fig. 2a, the qualitatively opposite behaviors manifest themselves in almost the entire region of temperatures and magnetic fields. Looking at the results near the transition lines [Fig. 2b], we find that the e-e interaction can enhance the WL correction when the SOC is weak.
Summary: We have formulated the renormalized GL theory including the Cooper and Coulomb interaction terms from the NLSM based on the Keldysh functional formalism. Our formalism extends the previous one and enables us to study dirty 2D Rashba superconductors under in-plane Zeeman fields, which are typical platforms of the SD effect. We find that the interplay between localization and interactions leads to the corrections that suppress the SC states, the first-order transition, and the SD effect. Furthermore, we have calculated the WL correction to conductivity that characterizes transport when the SC state is destroyed by an electric current exceeding the critical current. The WL correction can be suppressed or enhanced by interactions depending on the strength of SOC. We expect our results to pave the way for switching between superconducting, metallic, and insulating transport via an electric current.
Acknowledgements — We appreciate helpful discussions with Ryusuke Ikeda, Hideki Narita, Teruo Ono, and Ryuji Hakuno. This work was supported by JSPS KAKENHI (Grant Numbers JP22H01181, JP22H04933, JP23K17353, JP23K22452, JP24K21530, JP24H00007, JP25H01249).